OpenStudy (anonymous):

find -8 divided by -1/2. a)1/16 b)16 c)-1/4 d)-4

OpenStudy (anonymous):

thanks

OpenStudy (mathstudent55):

\[-8 \div (-\frac{ 1 }{ 2 })\] \(= -\dfrac{8}{1} \times (- \dfrac{2}{1})\) \(= + \dfrac{16}{1} \) = 16

OpenStudy (mathstudent55):

Answer is 16. Remember that to divide a fraction by a fraction, you multiply the first fraction by the reciprocal of the second fraction.

OpenStudy (anonymous):

your right i made a mistake

OpenStudy (anonymous):

sorry about that

OpenStudy (mathstudent55):

NP. I just hope the original poster looks in here again.

OpenStudy (anonymous):

lol you can go to her right now

OpenStudy (mathstudent55):

Good point. @gustina8 Look at the correct answer to this problem.

OpenStudy (anonymous):

got it,thanks!!